# Underdetermined Fourier Extensions for Surface Partial Differential Equations

### v2

@@ 3.7. Boundedness, Solvability, and Convergence
where we only include the $u^{\left(\tilde{N}\right)}\left(\boldsymbol{a}\right)=1$ condition in the case $f=g=0$ to ensure we get a non-zero solution. If there is a non-zero solution $u$ to the original PDE in $\mathcal{H}\left(d\right)$ (possibly requiring $u\left(\boldsymbol{a}\right)=1$ as well), then $u$ is feasible, and $\left\|u^{\left(\tilde{N}\right)}\right\|_{d}\leq\left\|u\right\|_{d}$ for all $\tilde{N}$. We have shown convergence of $\mathcal{F}u^{\left(\tilde{N}\right)}\to\mathcal{F}u$ in this case (Proposition 9). However, particularly for eigenvalue problems, we may also be interested in sufficient conditions for a solution to exist in $\mathcal{H}\left(d\right)$, and for $u^{\left(\tilde{N}\right)}$ to converge to a solution.
In Theorem 7.1 of the work by Chandrasekaran, Gorman, and Mhaska [chand18], the authors prove a result that, in our notation, would show $\mathcal{F}u\to f$ and $\mathcal{G}u\to g$ pointwise when a solution to (3.15) exists. Our next result adds to this for our approach by noting that mere boundedness of $u^{\left(\tilde{N}\right)}$ (which would be implied by feasibility) is enough to imply that a solution to (3.15) must exist, and that $u^{\left(\tilde{N}\right)}$ will converge to a solution $u$ in the $d$-norm. For choices of $d$ such that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, this implies uniform convergence of $u^{\left(\tilde{N}\right)}$ and its derivatives of up to order $p$ to $u$ over all of $\Omega$.
@@
First, note that for all $m>n$, $u^{\left(m\right)}$ is feasible for (3.16) on $S_{n},\partial S_{n}$, since $S_{n}\subset S_{m},\partial S_{n}\subset\partial S_{m}$. The fact that $u^{\left(n\right)}$ is the solution with minimum $d$-norm implies $\left\{\left\|u^{\left(n\right)}\right\|\right\}_{n}$ is a non-decreasing sequence. $\left\{\left\|u^{\left(n\right)}\right\|\right\}_{n}$ is bounded above by $\tilde{Q}$, and therefore $\left\|u^{\left(n\right)}\right\|_{d}\uparrow Q$ for some $Q>0$ ($u^{\left(n\right)}$ cannot be zero since either $u^{\left(n\right)}\left(\boldsymbol{a}\right)=1$ or one of $f,g$ is non-zero, and $u^{\left(n\right)}$ is $C^{p}$ since $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$ implies uniform convergence of the $u^{\left(n\right)}$ Fourier series and its derivatives of up to order $p$).
@@
$\left\|\partial^{\alpha}\left(u^{\left(n\right)}-u\right)\right\|_{L^{\infty}\left(\Omega\right)}$ $\leq\sum_{j=1}^{\infty}\left\|\boldsymbol{\omega}_{j}\right\|_{2}^{q}d^{-\frac{1}{2}}d^{\frac{1}{2}}\left|\hat{u}_{j}^{\left(n\right)}-\hat{u}_{j}\right|$
[…]
@@ 4.1. Surface Poisson
$S=\left\{\left(\cosh\left(t\right)\cos\left(s\right),\cosh\left(t\right)\sin\left(s\right),t\right):s\in\left[0,2\pi\right),\left(t-0.1\sin\left(3s\right)\right)\in\left[-1,1\right]\right\}.$
@@
The exact solution to this problem is $u\left(s,t\right)=\cos\left(4s\right)$. To solve this, we impose Hermite-Birkhoff interpolation conditions on a point cloud $S_{\tilde{N}}$ and boundary point cloud $\partial S_{\tilde{N}_{\partial}}$ to attempt to find a first-order Closest Point-like extension that approximately solves the PDE (see Proposition 3). Note that such an extension exists due to Proposition 5. Our discretized problem is:
@@
Our choice of $d_{n}$ is $\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2}$ with $q=4$; this choice achieves super-algebraic convergence when there is a solution in the native space. This is since the sequences $F^{\left(j,q\right)}d^{-\frac{1}{2}}$ and $G^{\left(j,q\right)}d^{-\frac{1}{2}}$ from Propositions 9 and 10, respectively, will both be in $\ell^{2}$ for all $q$ for this problem and choice of $d$. $T$ can be used to control the oscillation width of the $\psi$ functions from Subsection 3.3. We use $T=2$, which will provide fast convergence at the cost of poorer conditioning for dense point clouds, compared to smaller values of $T$. $T$ serves the same purpose as RBF shape parameters; however, conditioning here is not as much of an issue compared to most RBF methods, meaning a wider range of parameters are suitable. We discuss this further at the end of this subsection.
The other parameters used are $\ell=4$ for a $4\times 4\times 4$ extension domain $\Omega$, $N_{b}=\left(2\cdot 13+1\right)^{3}$ Fourier basis functions, and $\tilde{N}_{s}$ unique $s$ values so that $h_{\text{max}}\propto\tilde{N}_{s}^{-1}$. A plot of the computed solution for $\tilde{N}_{s}=100$ is in Figure 1 and a convergence test recorded in Table 1.
@@ 4.2. Eigenvalue Problem
$-\Delta_{S}u-\lambda u=0\text{, for some $u\neq 0$}.$
@@
(4.2) $\left(-\Delta\tilde{u}-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{u}\right)\hat{\boldsymbol{n}}_{S}-\lambda\right)\biggr{|}_{S_{\tilde{N}}}=0,\quad\hat{\boldsymbol{n}}_{S}\cdot\nabla\tilde{u}\biggr{|}_{S_{\tilde{N}}}=0,\quad\tilde{u}\left(0,0,1\right)=1.$
@@
The final condition ensures we produce a non-zero solution. At first glance, the location of the single point $\boldsymbol{a}$ where we enforce $\tilde{u}\left(\boldsymbol{a}\right)=1$ may seem to be important, since the eigenspace for a single eigenvalue may have $\tilde{u}\left(\boldsymbol{a}\right)=0$ for our selected $\boldsymbol{a}$. However, a random point would work with probability one, since the zero set of any eigenfunction is measure zero. That is, for arbitrary surfaces, choosing a point randomly should not cause an issue in practice. In this sphere example, symmetry ensures that there is an eigenfunction with $\tilde{u}\left(0,0,1\right)=1$ for any eigenvalue of $-\Delta_{S}$, and any point location would work. Another condition that forces the function to be non-zero would also be acceptable, such as setting a sum of function values or derivatives at random points to be equal to 1, but in our testing, setting $\tilde{u}\left(\boldsymbol{a}\right)=1$ for a single point $\boldsymbol{a}$ is sufficient. Eigenvalue multiplicities can be explored by setting multiple points where the function must be non-zero, but we leave this for later work.
@@
Table 2 again shows high-order convergence. Overall, this method of finding eigenvalues seems to be more reliable and theoretically justified than attempting to produce a differentiation matrix from interpolation followed by differentiation of the basis functions, which can produce incorrect results, even for small $\lambda$ for a variety of collocation methods (see the discussion at the end of Subsection 2.2 in [yan23]). Its use is also not limited to eigenvalues; any solvability problem relying on parameters can be investigated similarly.
@@ 5. Conclusions
We successfully developed, analyzed, and tested methods for Hermite-Birkhoff interpolation and PDEs on surfaces. The methods rely on underdetermined Fourier extensions, and work on unstructured point clouds with very few conditions for convergence. We have shown that our method produces solutions that approximately solve PDEs of interest in an extremely general setup (Propositions 9 and 10); convergence rates could then be estimated for a wide range of PDEs. Convergence is also shown in a general setting without rates by Proposition 12.
Of particular interest, our method works for surface PDEs on unstructured point clouds and is able to achieve arbitrarily high rates of convergence. For bulk problems, our method could be seen as an alteration of earlier, symmetric Hermite RBF methods (see, for example, [sun94, frank98, liu23]). Hermite RBFs do not seem to be commonly applied to surface PDEs outside of one approach for RBF finite differences [shaw19]. The setup of traditional Hermite RBF approaches, however, can be quite difficult, and may help explain why such methods are not widespread compared to non-symmetric RBF methods. We also find that it is typically easier to work with Fourier series directly to set up the correct linear system rather than with the often unwieldy functions that arise with Hermite RBFs.
@@
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN 2022-03302].
[…]

### v3

@@ 3.7. Boundedness, Solvability, and Convergence
where we only include the $u^{\left(\tilde{N}\right)}\left(\boldsymbol{a}\right)=1$ condition in the case $f=g=0$ to ensure we get a non-zero solution. If there is a non-zero solution $u$ to the original PDE in $\mathcal{H}\left(d\right)$ (possibly requiring $u\left(\boldsymbol{a}\right)=1$ as well), then $u$ is feasible, and $\left\|u^{\left(\tilde{N}\right)}\right\|_{d}\leq\left\|u\right\|_{d}$ for all $\tilde{N}$. We have shown convergence of $\mathcal{F}u^{\left(\tilde{N}\right)}\to\mathcal{F}u$ in this case (Proposition 8). However, particularly for eigenvalue problems, we may also be interested in sufficient conditions for a solution to exist in $\mathcal{H}\left(d\right)$, and for $u^{\left(\tilde{N}\right)}$ to converge to a solution. Note that if $\mathcal{F}$ can be written $\mathcal{F}^{\left(1\right)}+c\mathcal{F}^{\left(2\right)}$ for some function $c$, and we instead impose $\mathcal{F}^{\left(1\right)}=f$ and $\mathcal{F}^{\left(2\right)}=0$ on $S_{\tilde{N}}$, the following analysis does not change; this is typically how we write the Laplace-Beltrami operator, using Corollary 3.
In Theorem 7.1 of the work by Chandrasekaran et al. (2018), the authors prove a result that, in our notation, would show $\mathcal{F}u\to f$ and $\mathcal{G}u\to g$ pointwise when a solution to (3.9) exists. Our next result adds to this for our approach by noting that mere boundedness of $u^{\left(\tilde{N}\right)}$ (which would be implied by feasibility) is enough to imply that a solution to (3.9) must exist, and that $u^{\left(\tilde{N}\right)}$ will converge to a solution $u$ in the $d$-norm. For choices of $d$ such that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, this implies uniform convergence of $u^{\left(\tilde{N}\right)}$ and its derivatives of up to order $p$ to $u$ over all of $\Omega$.
@@
First, note that for all $m>n$, $u^{\left(m\right)}$ is feasible for (3.10) on $S_{n},\partial S_{n}$, since $S_{n}\subset S_{m},\partial S_{n}\subset\partial S_{m}$. The fact that $u^{\left(n\right)}$ is the solution with minimum $d$-norm implies $\left\{\left\|u^{\left(n\right)}\right\|_{d}\right\}_{n}$ is a non-decreasing sequence. $\left\{\left\|u^{\left(n\right)}\right\|_{d}\right\}_{n}$ is bounded above by $\tilde{Q}$, and therefore $\left\|u^{\left(n\right)}\right\|_{d}\uparrow Q$ for some $Q>0$ ($u^{\left(n\right)}$ cannot be zero since either $u^{\left(n\right)}\left(\boldsymbol{a}\right)=1$ or one of $f,g$ is non-zero, and $u^{\left(n\right)}$ is $C^{p}$ since $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$ implies uniform convergence of the $u^{\left(n\right)}$ Fourier series and its derivatives of up to order $p$).
@@
$\left\|\partial^{\alpha}\left(u^{\left(n\right)}-u\right)\right\|_{L^{\infty}\left(\Omega\right)}$ $\leq\sum_{j=1}^{\infty}\left\|\boldsymbol{\omega}_{j}\right\|_{2}^{q}d_{j}^{-\frac{1}{2}}d_{j}^{\frac{1}{2}}\left|\hat{u}_{j}^{\left(n\right)}-\hat{u}_{j}\right|$
$\leq\left\|\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\right\|_{\ell^{2}}\left\|u^{\left(n\right)}-u\right\|_{d}\text{, by Cauchy-Schwarz}$
@@
Proposition 10 shows convergence of $u^{\left(n\right)}$ to a solution of the PDE (3.9) without a rate estimate. To obtain a rate estimate, we need our PDE to have a unique solution and satisfy certain stability or regularity properties. In this case, Proposition 8 provides a straightforward way to prove high-order convergence of $u^{\left(n\right)}\to u$. We prove convergence here for operators satisfying a regularity condition; this condition is satisfied by elliptic operators under certain conditions.
Proposition 11.
Suppose $S$ is bounded and that for all $f\in C^{q}\left(\overline{S}\right),g\in C^{q}\left(\partial S\right)$ (if $S$ has boundary) there is a unique solution $u\in H^{r}(S)$ to (3.9) for some $r>0$ that can be extended to $\mathcal{H}\left(d\right)$. Also, assume that a regularity condition holds for (3.9) such that there exist constants $A,B>0$ and $\tilde{p}\geq 0$ such that for each $v\in H^{r}\left(S\right)$
(3.13) $\left\|v\right\|_{H^{r}\left(S\right)}\leq A\left\|\mathcal{F}v\right\|_{L^{2}\left(S\right)}+B\left\|\mathcal{G}v\right\|_{H^{\tilde{p}}\left(\partial S\right)}$
Then, let $u^{\left(\tilde{N}\right)}$ be the solution to (3.10) without the $u^{\left(\tilde{N}\right)}\left(a\right)=1$ condition. Assume $S$ and $\partial S$ satisfy the assumptions for $S^{\left(j\right)}$ in Proposition (8). Then, for small enough $h_{\text{max}}$, there are constants $A_{q},B_{q}>0$ such that
$\left\|u^{\left(\tilde{N}\right)}-u\right\|_{H^{r}\left(S\right)}$ $\leq A_{q}h_{\text{max}}^{q-\frac{m_{S}}{2}}\left\|u\right\|_{d}+{B}_{q}h_{\text{max}}^{q-\frac{m_{S}-1}{2}-\lceil{\tilde{p}}\rceil}\left\|u\right\|_{d}.$
Proof.
Assumption (3.13) applied to $u^{\left(\tilde{N}\right)}-u$ along with Proposition 8. ∎
There are a few comments to make regarding Proposition 11. First, note that since Proposition 8 gives global estimates, not just estimates on a finite set of points as in a finite difference consistency result, the regularity or stability of the PDE (3.9) itself is sufficient for convergence of the numerical method. We also point out that some conditions for the domain are implicitly imposed by the assumption that the solution $u$ to (3.9) can be extended to $\mathcal{H}\left(d\right)$. For Lipschitz flat domains $S$, functions on $H^{k}\left(S\right)$ can be extended to $H^{k}\left(\mathbb{R}^{m}\right)$ (see McLean, 2000, Thm. A.4). For manifolds with boundary, this theorem must be applied on patches to suitably extend $u$ to a larger manifold using a partition of unity, and then Proposition 4 can be applied. We can construct an $H^{k}$ periodic function on the box $\Omega$ by multiplying the extension to $\mathbb{R}^{m}$ by suitable $C^{\infty}$ bump functions (see Lee, 2003, Prop. 2.25). Finally, we comment that the regularity assumption (3.13) holds for elliptic problems on flat domains with unique solutions for Neumann, Dirichlet, or Robin boundary conditions, with various values of $r,\tilde{p}$ depending on the degree of smoothness of $\partial S$ and the coefficient functions of $\mathcal{F}$ (see McLean, 2000; Evans, 2010, Thms. 4.10, 4.11 and Thm. 4 of 6.3.2, respectively). For the regularity of the Laplace-Beltrami problem on a closed surface, see Thm. 3.3 of Dziuk and Elliott (2013), and for boundary value problems involving the Laplace-Beltrami operator on manifolds, see Prop 1.2, Thm 1.3, Prop 1.7, Eq. (7.6), and Prop. 7.5 in Chapter 5 of Taylor (2023).
@@ 4.1. Surface Poisson
(4.1) $S=\left\{\left(\cosh\left(t\right)\cos\left(s\right),\cosh\left(t\right)\sin\left(s\right),t\right):s\in\left[0,2\pi\right),\left(t-0.1\sin\left(3s\right)\right)\in\left[-1,1\right]\right\}.$
@@
The exact solution to this problem is $u\left(s,t\right)=\cos\left(4s\right)$. To solve this, we impose Hermite-Birkhoff interpolation conditions on a point cloud $S_{\tilde{N}}$ and boundary point cloud $\partial S_{\tilde{N}_{\partial}}$ to attempt to find a first-order Closest Point-like extension that approximately solves the PDE (see Corollary 3). Note that such an extension exists due to Proposition 4. Our discretized problem is:
@@
Our choice of $d_{n}$ is $\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2}$ with $q=4$; this choice achieves super-algebraic convergence when there is a solution in the native space. This is since the sequences $F^{\left(j,q\right)}d^{-\frac{1}{2}}$ from Proposition 8 will be in $\ell^{2}$ for all $q$ for this problem and choice of $d$. $T$ can be used to control the oscillation width of the $\psi$ functions from Subsection 3.3. We use $T=2$, which will provide fast convergence at the cost of poorer conditioning for dense point clouds, compared to smaller values of $T$. $T$ serves the same purpose as RBF shape parameters; however, conditioning here is not as much of an issue compared to most RBF methods, meaning a wider range of parameters are suitable. We discuss this further at the end of this subsection. The other parameters used are $\ell=4$ for a $4\times 4\times 4$ extension domain $\Omega$ and $N_{b}=27^{3}$ Fourier basis functions. This choice for $N_{b}$ ensures that the error is primarily determined by $h_{\text{max}}$ rather than the truncation of the Fourier series; $N_{b}\gg\tilde{N}$ for this test.
@@
We place points on the surface by constructing a tensor grid of $\tilde{N}=\tilde{N}_{s}^{2}/2$ points on the plane of $\left(s,t\right)$ coordinates in $\left[0,2\pi\right)\times\left[-1,1\right]$, where $\tilde{N}_{s}$ is the number of unique $s$-coordinate values so that $h_{\text{max}}\propto\tilde{N}_{s}^{-1}$. Then, we shift the values of $t$ up by $0.1\sin\left(3s\right)$ and map the points to $S$ with the parametrization given by (4.1). Importantly, the choice of points is not prescribed by the method, and we can choose our points freely as long as $h_{\text{max}}\to 0$ as $\tilde{N}\to\infty$. For example, a randomly spaced point cloud is used in the next subsection. However, more regularly spaced point distributions tend to produce lower errors in practice and should be used when possible. Figure 1 shows a plot of the computed solution for $\tilde{N}_{s}=100$ and Table 1 presents a convergence test.
@@ 4.2. Eigenvalue Problem
$\Delta_{S}u-\lambda u=0\text{, for some $u\neq 0$}.$
@@
(4.3) $\left(-\Delta\tilde{u}-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{u}\right)\hat{\boldsymbol{n}}_{S}-\lambda\tilde{u}\right)\biggr{|}_{S_{\tilde{N}}}=0,\quad\hat{\boldsymbol{n}}_{S}\cdot\nabla\tilde{u}\biggr{|}_{S_{\tilde{N}}}=0,\quad\tilde{u}\left(0,0,1\right)=1.$
@@
The final condition ensures that we produce a non-zero solution. At first glance, the location of the single point $\boldsymbol{a}$ where we enforce $\tilde{u}\left(\boldsymbol{a}\right)=1$ may seem to be important, since the eigenspace for a single eigenvalue may have $\tilde{u}\left(\boldsymbol{a}\right)=0$ for our selected $\boldsymbol{a}$. However, a random point would work with probability one, since the zero set of any eigenfunction is measure zero. That is, for arbitrary surfaces, choosing a point randomly should not cause an issue in practice. In this sphere example, symmetry ensures that there is an eigenfunction with $\tilde{u}\left(0,0,1\right)=1$ for any eigenvalue of $\Delta_{S}$, and any point location would work. Another condition that forces the function to be non-zero would also be acceptable, such as setting a sum of function values or derivatives at random points to be equal to 1, but in our testing, setting $\tilde{u}\left(\boldsymbol{a}\right)=1$ for a single point $\boldsymbol{a}$ is sufficient. Eigenvalue multiplicities can be explored by setting multiple points where the function must be non-zero, but we leave this for later work.
@@
Table 2 again shows high-order convergence. Overall, this method of finding eigenvalues seems to be more reliable and theoretically justified than attempting to produce a differentiation matrix from interpolation followed by differentiation of the basis functions, which can produce incorrect results, even for small $\lambda$ for a variety of collocation methods (see the discussion at the end of Subsection 2.2 from Yan et al. (2023)). Its use is also not limited to eigenvalues; any solvability problem relying on parameters can be investigated similarly.
@@ 5. Conclusions
We successfully developed, analyzed, and tested methods for Hermite-Birkhoff interpolation and PDEs on surfaces. The methods rely on underdetermined Fourier extensions and work on unstructured point clouds with very few conditions for convergence. We have shown that our method produces solutions that approximately solve PDEs of interest in an extremely general setup (Proposition 8); convergence rates could then be estimated for a wide range of PDEs, including those that satisfy a regularity condition (Proposition 11). Convergence is also shown in a general setting without rates by Proposition 10.
Of particular interest, our method works for surface PDEs on unstructured point clouds and is able to achieve arbitrarily high rates of convergence. For bulk problems, our method could be seen as an alteration of earlier, symmetric Hermite RBF methods (see, for example, Sun, 1994; Franke and Schaback, 1998; Liu et al., 2023). Hermite RBFs do not seem to be commonly applied to surface PDEs outside of one approach for RBF finite differences from Shaw (2019). The setup of traditional Hermite RBF approaches, however, can be quite difficult, which may help explain why such methods are not widespread compared to non-symmetric RBF methods. We also find that it is typically easier to work with Fourier series directly to set up the correct linear system rather than with the often unwieldy functions that arise with Hermite RBFs.
@@
Funding
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference numbers 579365-2023 (DV), RGPIN-2022-03302 (SR)].